OpenStudy (anonymous):

Let the set A = {3,5,6} and let the function f:A→A and g:A→A be defined by f(3)=5 , f(5) = 3 , f(6) =3 and g(3) = 3 , g(5) = 6 ,g(6) = 5 . Then compute the product function gof and fog.

OpenStudy (anonymous):

can anyone out there please help me ? .. like real fast ?

OpenStudy (anonymous):

gof can be defined as : gof(3)=g(5)= 6.... gof (5)=3...gof(6)=3

OpenStudy (anonymous):

can you solve it fully .. because i dont understand it at all !.. sorry :/

OpenStudy (anonymous):

okk it's easy...gof(3)=g(f(3))...and f(3)=5....gof(3)=g(5)=6... I hope you will understand !

OpenStudy (anonymous):

is that the full answer ?

OpenStudy (anonymous):

no ! you can do the others by yourself...gof(5)=g(f(5))...f(5)=...